Solve the second order differential equation (you can use any method) y + y = csc(x)
Answer #1

Q. Solve y+y= (se (x) Sol Iwe will solve this eě by variation of parameter. y+y = Csc (2) , R(x) = (SC(x) Csc (2) (6²+1) 9le Cosxx Coca = f loszx dx Cotx dx 11 بع rea Log I sin (D) Thus lp. I -X Cos x xt sinx Sinz (loglsinx)) General sol C.F. + P